OpenStudy (anonymous):

There are 4 dogs. Aslan, rover,spot, and prince. How many fleas does aslan have? Rover & spot have 35 fleas in between them Rover has two thirds as many fleas as aslan Prince and spot have 38 fleas altogether rover and prince have 39 fleas altogether

OpenStudy (anonymous):

27

OpenStudy (anonymous):

exactly how did you get to that?

OpenStudy (anonymous):

rover + spot = 35 spot + prince = 38 so prince = rover +3 prince + rover = 39 rover +3 + rover = 39 rover = 18 rover = 2/3 aslan aslan = rover * 3/2 = 18*3/2 = 27

OpenStudy (anonymous):

wow thanz you guys are the best!

OpenStudy (anonymous):

well if R+S=35 and P+S=38 and R+P=39 then R+S+S+P=35+38 and if R+P=39 then (R+S+S+P)-(R+P)=(35+38)-(39) so S+S=34 S=17 if P+S=38 then P=21 if R+S=35 then R=18 so S=17 R=18 and P=21 if R=2/3A then 18=2/3A so 9=1/3A thus 27=A

OpenStudy (anonymous):

thats another way... XD

OpenStudy (anonymous):

thanks so much arman1002!!!!